### v1

# Overcoming False Illusions in Real-World Face Restoration with Multi-Modal Guided Diffusion Model

Abstract
We introduce a novel Multi-modal Guided Real-World Face Restoration (MGFR) technique designed to improve the quality of facial image restoration from low-quality inputs. Leveraging a blend of attribute text prompts, high-quality reference images, and identity information, MGFR can mitigate the generation of false facial attributes and identities often associated with generative face restoration methods. By incorporating a dual-control adapter and a two-stage training strategy, our method effectively utilizes multi-modal prior information for targeted restoration tasks. We also present the Reface-HQ dataset, comprising over 23,000 high-resolution facial images across 5,000 identities, to address the need for reference face training images. Our approach achieves superior visual quality in restoring facial details under severe degradation and allows for controlled restoration processes, enhancing the accuracy of identity preservation and attribute correction. Including negative quality samples and attribute prompts in the training further refines the model’s ability to generate detailed and perceptually accurate images.
Figure 1: The proposed MGFR model demonstrates an exceptional ability in restoring low-quality face images, yielding more outstanding visual effects with the addition of reference images, particularly in situations of extreme degradation, shown in (a). Furthermore, the model is capable of target-specific restoration in (b), directed by facial attribute prompts. This encompasses defining facial age characteristics (Case 1), adjusting the restoration process based on attribute prompts (Case 2), and executing precise modifications to facial elements (Case 3). w/o Reference Image means the results of our model without introducing reference image.
1 Introduction
Real-World Face Restoration (FR) aims to reconstruct high-resolution, high-quality (HQ) facial images from their degraded, low-resolution observations. Recent works, leveraging powerful generative priors and diffusion models, have achieved significant progress Menon et al. (2020); Yang et al. (2021b); Wang et al. (2021b); Lin et al. (2023); Wang et al. (2023b), particularly in addressing severely degraded facial images. However, the information contained in the low-quality (LQ) inputs is limited. FR inevitably introduces the illusion of generation, producing results with different facial attributes or even different identities from the target image. For example, in Figure 1 (a) and Figure 2, we cannot effectively predict the eye colour and skin characteristics of the person in the LQ input, resulting in the output results – even the quality can be improved – having an apparent perceptual distance from the target image. Many applications find this unacceptable, as humans can readily identify these flaws. Achieving optimal facial image recovery requires effectively tackling false hallucinations.
Practically, we find that for the restoration of specific face images, we can obtain a lot of prior information. For example, we may know this person’s various attributes and identity, and there may even be other clear images of this person in the photo album. Suppose we can use this information as additional guidance to guide the restoration. In that case, we can alleviate the impact of false illusions on key issues, thus helping to generate facial details that better suit our needs. For example, in Figure 2 (a), when we provide an additional key description of gender and age, we can correct the illusion. In Figure 1 (a) and Figure 2 (b), additional high-quality images are used as reference, and the details of the eyes and skin texture can be accurately generated. What is even more gratifying is that this kind of prior information can be widely obtained, making this problem of application significant.
This work proposes a method called Multi-modal Guided Real-World Face Restoration (MGFR). We aim to use multiple control methods to consider diverse multi-modal prior information in FR to restore face images in a targeted manner. Specifically, MGFR uses attribute text prompts, HQ reference images, and identity information as priors for collaborative guidance during restoration. We design a dual-control adapter with a two-stage training strategy to balance the complex multi-modal and multi-source prior information. This dual controller is compatible with pre-trained generative diffusion models Rombach et al. (2022) and prioritises restoration tasks while incorporating additional multi-modal guidance. In addition, we collect the Reface-HQ dataset to address the scarcity of reference image samples containing over 5000 identities and 23000 high-resolution facial images. Based on the FFHQ Karras et al. (2019a) and the proposed Reface-HQ datasets, we develop a high-quality synthetic dataset for model training enriched with attribute text prompts. Furthermore, we adopt a counter-intuitive strategy to integrate negative-quality samples with negative-quality prompts and negative-attribute prompts into training to enhance perceptual quality and detail generation.
The proposed MGFR model shows exemplary performance in the FR task, achieving superior visual quality in facial details, especially under severe degradation conditions. MGFR can take a high-resolution reference image as prior information and restore important details based on the reference image that cannot be displayed in the LQ input. The identity information provided by the reference image will also be considered in FR to ensure that the restoration does not change the identity characteristics. In addition, MGFR can also provide a certain degree of control over the restoration process through attribute text prompts, significantly enhancing the feasibility of FR. As shown in Figure 1 (b), textual prompts fulfil a dual function: they significantly reduce facial attribute illusions, such as “big eyes” or “old”, and also guide the restoration of specific facial features, such as “wearing glasses” and “young”.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption (a)     LQ No Prompt ‘woman, young, no beard…’ GT (b)     LQ w/o Reference w/ Reference GT
Figure 2: Motivation. In conditions of severe degradation, the loss of facial identity information becomes profoundly pronounced without reference image. During the face restoration process, distortions of facial attributes, including gender and age, are commonly encountered. Appropriate attribute prompts can offer additional reference points and exert control in the recovery process.
2 Related Works
Real-World Face Image Restoration
Real-World face restoration (FR) concentrates on the challenging task of reconstructing HQ face images from LQ inputs. These LQ inputs are often blemished by various forms of quality degradation such as low-resolution Chen et al. (2018); Dong et al. (2014); Lim et al. (2017), blur Kupyn et al. (2018); Shen et al. (2018), noise Zhang et al. (2017), and JPEG compression artifacts Dong et al. (2015), etc. FR heavily relies on facial priors, such as facial landmarks Chen et al. (2018), parsing maps Chen et al. (2018; 2021), and facial component heatmaps Yu et al. (2018). Generative priors Karras et al. (2020); Rombach et al. (2022); Gu et al. (2020); Shen et al. (2020) have also emerged as fundamental elements in providing vibrant textures and details in FR Menon et al. (2020); Hu et al. (2023); Zhu et al. (2022). Advanced techniques like GPEN Yang et al. (2021a), GFP-GAN Wang et al. (2021a), and GLEAN Chan et al. (2021) are recognized for more effectively incorporating these priors within encoder-decoder structures. There are also works that considerably reduce the uncertainty commonly associated with generative priors Gu et al. (2022); Zhou et al. (2022); Wang et al. (2022), which are trained on discrete feature codebooks for high-quality facial images. Recently, diffusion models like DiffBIR Lin et al. (2023) have revitalized interest in this area, leveraging the generative power of pre-trained LDM as a prior. DR2 Wang et al. (2023b) also makes contributions by transforming input images into noisy states and then iteratively denoising them to capture the essential semantic information.
Reference-Based Face Image Restoration
Reference-based face restoration utilizes HQ images of the same identity as references. This concept was first introduced in Li et al. (2018a). To address discrepancies in poses and expressions, GWAInet Dogan et al. (2019) and the later work of Li et al. Li et al. (2020; 2018b) focused on more effectively directing deformations or choosing the optimal reference image for reconstruction. MyStyle Nitzan et al. (2022) adopts a unique approach by refining StyleGAN Karras et al. (2019a) with numerous reference images based on personal appearance. DMDNet Li et al. (2023) employs a dictionary constructed from diverse, high-quality facial images to rehabilitate degraded images using its high-quality components. In the MGFR framework, incorporating a single reference image is vital for tailoring the restoration to individual faces. Unlike conventional methods, MGFR does not require strict alignment constraints on expressions or postures.
Multi-modal Guided Generation
Diffusion models have shown significant effectiveness in a broad range of image processing tasks. Current methods Chen et al. (2023b); Zhang et al. ; Yu et al. (2024); Chen et al. (2023a) employ pre-trained text-to-image diffusion models Rombach et al. (2022) for image processing, demonstrating the potential of language as a comprehensive input for image reconstruction tasks. Concurrently, approaches like ControlNet Zhang et al. , T2I-adapter Mou et al. (2023), and ControlNet-XS Zavadski et al. (2023) have further developed the integration of more intricate condition controls within the text-to-image framework, facilitating more precise and tailored image generation. Nevertheless, the field of FR, particularly in the utilization of natural language prompts, continues to be an area of untapped potential.
3 Methodology
The proposed MGFR method is able to take face attribute text prompts, reference images, and identity information as input to alleviate illusions and improve visual effects. MGFR involves controlling information from multiple modalities. However, we found that if the model is directly trained to process control information from multiple sources and of different importance, it is not easy to utilize all the information effectively. The model may ignore the more complex information to utilize. This causes some of the controls to fail and reduces image quality. In our method, text prompts are the most complex control information because they involve understanding text and the correspondence between text and face attributes. Therefore, we divide the training into two stages. In the first stage, the training focuses on the basic text-guided restoration model (Section 3.1). This allows the model to restore high-quality images and understand facial attributes. Then, we introduce other control information on this basis. The second stage introduces the HQ reference image and face identity information as the control means (Section 3.2). To improve the image effect further, Section 3.3 describes negative examples and the adopted prompting strategy.
Figure 3: Framework Overview. This figure illustrates the overall workflow of the proposed MGFR model.
3.1 Stage One: Text-guided Face Restoration
This stage trains the face image restoration model that accepts text prompt as input, as shown in Figure 5. We use the pre-trained Stable Diffusion (SD) Rombach et al. (2022) model as our generative prior and train an additional adapter to extend it to the face restoration applications. The pre-trained SD generative prior has the ability to understand face image attributes text and generate high-quality face images. In this stage, our model restores the image $x$ according to the condition $\{y,c_{a}\}$, where $y$ represents the degraded LQ image, and $c_{a}$ constitutes the facial attribute prompts describing the face attributes. We first use the CLIP text encoder Radford et al. (2021) to calculate the text embedding $e_{r}=\operatorname{CLIP}(c_{a})$. The LQ input $y$ is also mapped to a latent representation $z_{lq}$ using the VAE encoder in SD Rombach et al. (2022). We then perform diffusion generation on this latent representation. In the framework of SD, the model uses UNet Ronneberger et al. (2015) denoising model ${\mathcal{E}}_{\theta}\left({z}_{t},t,e_{r}\right)$ to perform the diffusion generation process, where $t$ is the time stamp in diffusion model and $z_{t}$ is the intermediate results at time $t$. Based on the ControlNet Zhang et al. framework, we use an external adapter that takes the LQ input $y$ and text prompts embedding $e_{r}$ as input to provide guidance for the fixed UNet $\mathcal{E}_{\theta}$. We call this adapter the LQ Control Adapter (LCA). Specifically, the UNet model contains the encoder, intermediate blocks, and the decoder. The decoder receives features from the encoder and fuses them at each corresponding scale. The LCA contains the same encoder and intermediate blocks as in the UNet model. The feature output of each scale in LCA is integrated with the corresponding scale of the UNet decoder to achieve the effect of output control. However, we found that simply using the above ControlNet framework has a key limitation – the lack of information exchange from the UNet encoder to the LCA. This gap means that the LCA is unaware of the processes that are performed in the UNet encoder, thereby limiting its ability to generate effective control features. In order to solve this problem, we add the feature output of each scale in the UNet encoder to the corresponding scale in the LCA. The LQ controller part of Figure 5 illustrates this operation. In this way, the capability of the LCA is greatly enhanced, so better visual effects and control results can be achieved.
3.2 Stage Two: Multi-modal-guided Face Restoration
After the first stage of training, the model can already reconstruct high-quality images from the LQ inputs, guided by text prompts. Next, we further enrich the guidance and introduce high-quality reference images and face identity information as additional control means based on the first-stage model. We design a new Dual-Control Adapter (DCA), as shown in Figure 3. In DCA, we introduce a Reference Control Adapter (RCA) specifically for reference image processing. RCA has the same architecture as LCA, and its role is to extract related and useful information from reference images and identity information and provide additional details to LCA. The input of RCA includes an HQ reference face image $r$ containing the same identity as the LQ input and its identity information embedding $e_{f}$. For the reference image $r$, we first use the VAE encoder consistent with the SD model for feature extraction to obtain $z_{ref}$. Next, we fuse the LQ latent representation $z_{lq}$ with $z_{ref}$ using a reference and LQ feature fusion module (RLF). This module allows RCA to identify the high-frequency details missing in the LQ input and perform targeted information extraction for restoration guidance. For identity information embedding, we calculate $e_{f}=\operatorname{Proj}(\operatorname{Arcface}(r))$, where $\operatorname{Arcface}(r)$ is the face recognition model Arcface Deng et al. (2019) to extract the identity feature from the reference image $r$. We align it to the space that RCA can handle through a trainable linear projection layer $\operatorname{Proj}(\cdot)$. Due to the function of the RCA extracting information from the reference image according to the LQ input, the RCA requires the information of the LCA branch as input. At the same time, RCA needs to provide the extracted information back to LCA in reverse. Therefore, we designed a dual-way interaction mechanism for RCA and LCA, as shown in Figure 5. In this design, RCA runs in parallel with LCA. At each scale, the LQ block in LCA first processes the fused information of both two branches and then hands the intermediate features to RCA. RCA performs feature extraction and processing based on these intermediate results and reference conditions and finally uses the same operation to apply the processing results to the next layer of LCA processing. Finally, the output of each scale of LCA is applied to the corresponding position of the UNet decoder. RCA directly affects LCA and, therefore, also affects the calculation of UNet. Since then, we have had a dual-control adapter that can accept multiple control inputs.
Figure 4: The model architecture employed during the initial training stage is discussed. In the article, ’Ours w/o Reference Image’ refers to the outcome of the model trained following this stage.
Figure 5: Dual-control Adapter. LQ-blocks are from the LQ control adapter (LCA), and Ref.-blocks are from the reference control adapter (RCA). $\oplus$ represents the element-wise add operation.
3.3 Negative Samples and Prompt
Classifier-Free Guidance (CFG) Ho & Salimans (2022) introduces a novel control mechanism utilizing negative prompts to delineate unwanted content for the model. This feature can be leveraged to inhibit the generation of low-quality images by the model and to enhance the precision of facial detail reconstruction. Throughout the inference phase, at each step of diffusion, three distinct predictions are generated: one employing the positive prompt $pos$, another using the negative quality prompt $nq$, and a third via the negative attribute prompt $na$ (the negation sentence described by $pos$). We combine the results generated from these different prompts to form the final output:
$\tilde{z}_{t-1}=z_{t-1}^{pos}+\lambda_{nq}\times(z_{t-1}^{pos}-z_{t-1}^{nq})+\lambda_{na}\times(z_{t-1}^{pos}-z_{t-1}^{na}),$ (1)
where $\lambda_{na}$ and $\lambda_{nd}$ is the hyperparameters, $z_{t-1}^{pos}=\mathcal{E}_{\theta}(z_{t},z_{lq},z_{ref},t,pos)$, $z_{t-1}^{nq}=\mathcal{E}_{\theta}(z_{t},z_{lq},z_{ref},t,nq)$, $z_{t-1}^{na}=\mathcal{E}_{\theta}(z_{t},z_{lq},z_{ref},t,na)$. $pos$ represents a standard description of a facial attribute. $nq$ is the negative words of quality, e.g., “oil painting, cartoon, blur, dirty, messy, low quality, deformation, low resolution, over-smooth”. $na$ is used for a negative description of a facial attribute, implying complete negation. Accurate prediction in both positive and negative directions is essential for the CFG technique. The lack of negative-quality samples and prompts in our training might cause the model to misinterpret negative prompts, leading to artifacts. To resolve this, we generated 16K images with negative-quality prompts using the original SD generative model and included these low-quality images in our training to enable the model to learn the concept of negative quality. Figure 9 (a) shows an example of the negative quality sample and prompt.
Figure 6: The MGFR model demonstrates a remarkable capacity for restoring LQ images. Upon integrating the reference image, particularly in instances of severe degradation, the model significantly enhances the restoration of facial details and overall image quality.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Input DiffBIR CodeFormer w/o Reference Ours Reference
Figure 7: In the qualitative comparison of real-world low-quality (LQ) images, MGFR demonstrates success in recovering facial details without false illusion and preserving identity from. Please zoom in for a better view.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Input No Prompt Add: ‘lipstick’ Input No Prompt Add: ‘earrings’ Input No Prompt Add:‘big eyes’
Figure 8: MGFR demonstrates capability of face image restoration facilitated by text prompts. It possesses the capacity to artificially modulate specific aspects of the restoration outcomes, such as determining the presence of accessories like lipstick or glasses (Cases 1 & 2), and orchestrating the restoration process in alignment with facial attributes (Case 3).
4 Experiments
4.1 Experimental Setting
Datasets. Our two-stage training method requires different datasets for training. For the first stage, we mainly train the model’s ability to restore HQ images and process text prompts. Therefore, we need HQ images with text annotations for training. We synthesize training image pairs using the FFHQ dataset Karras et al. (2019b). FFHQ contains 70,000 high-resolution face images, and we resize these images to $512\times 512$ for training. In the second stage, in addition to requiring HQ face images to create training image pairs, we also need to assign HQ reference images with consistent identities but different details to each image. Although there are some datasets proposed for reference face restoration Liu et al. (2015); Yi et al. (2014), the resolution and quality of these datasets cannot meet the current requirements. In this work, we collect a new dataset for referenced face restoration called Reface-HQ. Reface-HQ contains 23,500 high-quality and diverse images of over 5,250 identities. Additional details of Reface-HQ can be found in the Appendix A. To synthesize LQ images, we follow the degradation model and setting used in Wang et al. (2023b). Our test data also involves multiple sources, including CelebA-Test Liu et al. (2015), Reface-Test, and real-world LQ images collected from the Internet. Specifically, CelebA-Test contains 3,000 testing images from the CelebA-HQ dataset. Reface-Test contains 1,800 images of 380 identities split from the proposed Reface-HQ dataset. The LQ images for testing are synthesized within the same degradation range as the training setting.
Figure 9: Training Data Composition. Initially, negative quality samples are incorporated into the training to enhance the clarity and quality of the restored image. Furthermore, large language models coupled with a facial attribute classifier are employed to extract attribute texts for integration into the training.
Attribute Prompt. Text prompts are important for us to control face attributes and improve quality. In our method, a total of three types of prompts are introduced. Two attribute prompts describe the face attributes, and the last one describes the negative quality of the image. For attribute prompts, $pos$ contains positive descriptions of face attributes, while $na$ describes attributes that do not exist in this image to provide negative prompts of attributes. To obtain these descriptions, we first use a pre-trained face attribute detector He et al. (2017) to extract the presence of each attribute in the face. We considered 28 different attributes in this work. For attributes that have high confidence to exist, we add them to the $pos$ positive attributes. For the remaining attributes with low confidence, we classify them as $na$ negative attributes. At this time, these attributes are still separate words. We use a large language model to organize the separated words into natural language to facilitate the understanding of the CLIP text encoder. Thus, each face image is associated with two attribute prompts detailing existing and non-existing attributes. For the negative quality prompt, $nq$ involves “low quality, low resolution, over-smoothed and distorted images”, as shown in Figure 9 (a). See Appendix B for comprehensive details on the training and inference procedures involving attribute prompts.
Implementation. The training involved fine-tuning based on Stable Diffusion v2.1 Rombach et al. (2022), with the control adapter structure adhering to Zhang et al. . The Adam optimizer Kingma & Ba (2014) was employed, featuring a learning rate of $e^{-5}$. The initial training stage spanned 15 days, while the subsequent stage lasted 5 days, utilizing 4 Nvidia A100 GPUs with a batch size of 4. For testing purposes, the hyperparameters were set as $T=500$, $\lambda_{na}$ = 0.5 and $\lambda_{nq}$ = 0.5.
Metrics. For quantitative comparison, followed by many previous works Lin et al. (2023); Yu et al. (2024), the selected metrics include full-reference metrics PSNR, SSIM, and LPIPS Zhang et al. (2018), as well as non-reference metrics ManIQA Yang et al. (2022), ClipIQA Wang et al. (2023a), and MUSIQ Ke et al. (2021). Furthermore, the Arcface identity distance Deng et al. (2019) (ID) is utilized to assess the similarity of identity information.
Real-SR($\times$4) Real-SR($\times$8) Real-SR($\times$16)
Method LPIPS ↓ ManIQA ClipIQA MUSIQ LPIPS ↓ ManIQA ClipIQA MUSIQ LPIPS ↓ ManIQA ClipIQA MUSIQ
PSFRGAN 0.2938 0.5927 0.5702 73.39 0.3315 0.6015 0.5956 73.08 0.3788 0.5739 0.6274 71.76
GPEN 0.2828 0.6596 0.6430 69.25 0.3217 0.6754 0.6299 68.63 0.3831 0.6618 0.5897 66.61
VQFR 0.2951 0.2875 0.2490 62.95 0.3277 0.4163 0.2363 61.92 0.3761 0.6513 0.2148 60.49
CodeFormer 0.2927 0.5803 0.5179 75.47 0.3193 0.5970 0.6235 75.09 0.3821 0.5803 0.5877 70.85
DR2 0.3264 0.5749 0.4441 63.43 0.3580 0.5246 0.4494 59.46 0.3796 0.5160 0.5035 70.31
DiffBIR 0.2611 0.6068 0.7681 74.27 0.3017 0.6058 0.7439 73.87 0.4238 0.5361 0.7164 67.41
Ours w/o Reference 0.2925 0.6854 0.8244 76.22 0.3227 0.6776 0.8083 75.94 0.3760 0.6729 0.7944 75.76
Table 1: Quantitative Comparison in CelebA-Test. Results in red and blue signify the highest and second highest, respectively. The $\downarrow$ indicates metrics whereby lower values constitute improved outcomes, with higher values preferred for all other metrics.
4.2 Comparisons with State-of-the-art Methods
MGFR is qualitatively and quantitatively compared with state-of-the-art methods in FR. Notably, the model trained in the initial stage, which is a restoration model solely guided by attribute prompts, already achieves superior visual results. The non-reference prior-based methods selected include PSFRGAN Chen et al. (2021), GPEN Yang et al. (2021a), VQFR Gu et al. (2022), CodeFormer Zhou et al. (2022), DR2 Wang et al. (2023b), and DiffBIR Lin et al. (2023), along with reference prior-based methods ASFFNet Li et al. (2020) and DMDNet Li et al. (2023). Particularly, to ensure contrastive fairness during the inference stage, the description text, containing restricted attributes, is obtained through low-resolution processing. In practical applications, however, users can freely set attribute prompts, enabling more precise and comprehensive guidance. For qualitative results comparing ASFFNet and DMDNet, please refer to the Appendix.
Refer to caption Refer to caption Refer to caption Refer to caption Input No Prompt ‘No beard, young’ ‘young, woman blonde hair’
Figure 10: Attribute prompts that manifestly contravene low-resolution inputs prove ineffectual and result in distortions and artifacts within the restored image.
Refer to caption Refer to caption Refer to caption Refer to caption Input High Blur Face Swap Reference
Figure 11: Face Swapping: MGFR is capable of leveraging the reference map to alter the comprehensive components of the face.
Refer to caption Refer to caption (a) User Study on compared methods (b) User study on using Reference image
Figure 12: The results of our user study. We randomly select face images under multiple test datasets for user study. Our model achieves excellent recovery quality, which can be further enhanced with high quality reference image and identity information guidance.
Comparison on Synthetic Degradations. Firstly, a quantitative comparison of our model without reference images on the synthetically degraded CelebA-Test dataset is conducted without reference image guidance. According to Table 1, our model achieves the best results on all non-reference metrics, indicative of the superior image quality of the results. Due to space limitation, values of SSIM and PSNR of Table 1 are shown in Section C.1. Additionally, the method’s limitations on full-reference metrics are also noted. This phenomenon, preliminarily demonstrated by experiments in Yu et al. (2024); Jinjin et al. (2020), necessitates a reevaluation of the reference value of indicators like PSNR, SSIM, LPIPS, and the proposal of more effective methods to assess advanced FR methods, particularly as quality improves. More qualitative comparison results of our model can be found in Appendix C. Subsequently, Figure 6 and Figure 1 (a) present a qualitative comparison of the MGFR method applied to the Reface-Test dataset. Even in cases of severe degradation, our method successfully produces highly superior facial details guided by the reference image. In addition, we provide a comparison between our model without reference images and MGFR, with a particular focus on FR tasks involving features like double eyelids, pupil color, and finer facial details, such as wrinkles and moles, which cannot be accurately captured without reference image guidance. This further demonstrates the superiority of utilizing reference image guidance in the FR task. Finally, Table 2 offers quantitative comparison results, indicating that our method significantly surpasses other state-of-the-art methods in perceived quality.
We also conduct a user study with a total of 40 participants, comparing MGFR to other approaches. Participants were asked to select the best quality recovery result from these test techniques for each pair of comparison images, or if no reference image was provided, the result that came closest to the Ground Truth. Figure 12 presents the results, which demonstrate that our method outperforms the state-of-the-art methods in terms of recovery quality. Furthermore, the reconstruction effect can be further enhanced by using the reference image guidance.
Degradation Method PSNR SSIM LPIPS ↓ ManIQA ClipIQA MUSIQ ID ↓
ASFFNet 23.43 0.6811 0.2452 0.5685 0.6215 71.66 0.7053
DMDNet 23.85 0.7062 0.2667 0.5023 0.6023 72.31 0.6964
DR2 23.58 0.6581 0.2532 0.5340 0.5956 69.00 0.7957
CodeFormer 23.88 0.6904 0.2912 0.4959 0.5823 74.80 0.6579
DiffBIR 24.12 0.6717 0.2785 0.5547 0.7474 73.73 0.6379
$\times$8 MGBFR(Ours) 23.10 0.6248 0.2688 0.6535 0.8147 75.51 0.5166
ASFFNet 21.70 0.6472 0.3013 0.5803 0.6221 71.57 0.9361
DMDNet 22.37 0.6761 0.3179 0.4579 0.4727 67.27 0.9270
DR2 22.28 0.6720 0.3269 0.5233 0.5693 66.39 0.8676
CodeFormer 21.88 0.6124 0.3400 0.5547 0.5855 71.30 0.8658
DiffBIR 21.51 0.5939 0.3944 0.4937 0.7144 67.42 0.8876
$\times$16 MGBFR(Ours) 21.75 0.6033 0.2989 0.6524 0.8046 75.06 0.7401
Table 2: Quantitative Comparison in Reface-Test. Quantitative comparison of guided recovery results based on reference images. DR2, CodeFormer, and DiffBIR do not use reference images.
Comparison on Real-world Degradations. Additionally, our method was tested on real-world LQ images, which involved collecting degraded face images of publicly available images alongside reference images. The qualitative results, presented in Figure 12, demonstrate that the resulting images possess realistic visual effects with minimal facial illusions. More quantitative and qualitative results are presented in Appendix C.
4.3 Controlling Restoration with Attributes Prompts
Our method facilitates targeted image restoration guided by attribute prompts. As illustrated in Figure 8, the comparison between the first and second cases reveals that the integration of supplementary attribute prompts facilitates the manipulation of subtle facial attributes absent in the original image. This includes the addition of glasses, earrings, and accessories. In scenarios of severe degradation, exemplified by the third case, reconstructing facial features like eyes poses a significant challenge without external prompts. More results are shown in Section E.1.
However, it is imperative to acknowledge that attribute prompts do not invariably yield efficacy. As demonstrated in Figure 11, our model is capable of control tuning through attribute prompts. However, prompts that starkly contradict LQ inputs, like “blonde hair”, are found to be ineffective. This ensures the model’s adherence to the provided LQ inputs. Furthermore, as illustrated by the input of an LQ male face in Figure 11, when the input attribute is “Female”, the model subtly incorporates the attribute label “Female” into the image. This is achieved through modifications like the addition of an earring and the removal of the beard while remaining faithful to the LQ input. Such modifications further underscore the efficacy of attribute text in guiding the restoration process. This outcome is not unexpected. On the contrary, excessive control capability might lead to a reduction in the restoration effectiveness, countering the fundamental intent of image reconstruction efforts and thereby demonstrating the robustness of the proposed method.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Input w/o Negative Prompt Use Negative Quality Prompt Use Negative Attribute Prompt GT
Figure 13: Negative quality prompts engender restoration outcomes characterized by high definition, whereas negative attribute prompts yield results with enhanced detail.
4.4 Ablation Study
Real-SR(×8) SSIM PSNR LPIPS ManIQA ClipIQA MUSIQ
w/o Link-UL 0.6372 22.17 0.3275 0.5931 0.7082 71.33
w/o Link-LR 0.6659 23.86 0.2873 0.6152 0.6645 65.70
MGBFR(Ours) 0.6248 23.10 0.2688 0.6535 0.8147 75.51
Table 3: Ablation study of additional information exchange in the MGBFR model. ’w/o Link-LR’ means that the upward flow of information from LCA to RCA is removed.
Prompts
$pos$ $nq$ $na$ LPIPS ↓ SSIM PSNR ManIQA ClipIQA MUSIQ
0.3264 0.6858 25.15 0.4782 0.2568 49.97
0.2690 0.6484 24.43 0.6441 0.7008 73.26
0.2930 0.6066 23.27 0.6656 0.7999 75.34
0.2702 0.6511 24.49 0.6437 0.7029 73.11
0.3227 0.5904 22.34 0.6776 0.8083 75.94
Table 4: Ablation study of attribute prompts and negative prompts
Attribute Prompt and Negative Samples. Figure 13 displays qualitative results under various settings, aligning with the strategies outlined in Section 3.3. It can be seen that incorporating a negative quality prompt significantly enhances restoration quality, while the addition of a negative attribute prompt yields images with finer details. Quantitative results under various settings are also presented in Table 4. Adding either positive attribute prompts or negative quality prompts is observed to improve the perceived quality of the images significantly. Utilizing both types of prompts in conjunction with the negative attribute prompt achieves the most favourable perceived effect. The impact of hyperparameters on the results was also explored, revealing that settings of $\lambda_{na}=0.5$ and $\lambda_{nq}=0.5$ yield the best perceptual outcomes, balancing sharpness and definition. Please refer to Appendix G for detailed qualitative results in different hyperparameters.
Face Swapping. MGFR can facilitate face-swapping operations involving the processing of highly degraded LQ images to obscure identities, as shown in Figure 11. Face images and identity information from different identities are utilized as guides to achieve face swapping and identity replacement. Besides proposing an additional application for the model, this experiment further illustrates the method’s efficacy in utilizing identity information and reference images for guidance.
Additional Information Exchange. Unlike Zhang et al. , we have integrated an additional information flow exchange link (Link-UL) from the U-net model to LCA, and a bidirectional information flow link (Link-LR) between LCA and RCA. Table 4 displays the quantitative test results for the presence of the aforementioned information exchange links. Notably, ‘w/o Link-UL’ refers to results obtained with a single information flow from LCA to U-net model. It is evident that additional information flow exchanges result in improved perceived quality.
Arcface Identity Embedding. Our model is able to leverage identity information to guide the image restoration process, aiming to mitigate the deficit in facial identity information substantially. As demonstrated in Figure 15, our model employs the identity encoding formulated by the identity information extractor to mitigate the deficiency of facial identity information in the restored image. After losing arcface identity embedding, the recovered results still have high quality but there is a false illusion of face identity information.
Different expressions and poses reference images. In previous studies, reference image-based face restoration has been widely explored, but its efficacy is constrained by the need for strict alignment between the reference image and the low-quality (LQ) input. As shown in Figure 21, the recovery results of ASFFNet and MDMNet exhibit severe distortions when the reference image and the LQ input are slightly misaligned. However, our method completely resolves this issue, as it imposes no strict requirements on the expression, pose, or other variations of the reference image. As demonstrated in Figure 14, even when there are discrepancies between the reference image and the LQ input, such as face orientation, labeling, makeup, or pose, our model consistently achieves high-quality restoration without any artifacts.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption (a)     LQ Ours GT Reference (b)     LQ Ours GT Reference
Figure 14: Face restoration results with reference images with different expressions and poses. When there are differences in pose and expression between the reference image and the low-quality input image, our model can still achieve a good restoration effect without the generation of artifacts.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption (a)     LQ Ours w/o ID Ours w/ ID GT (b)     LQ Ours w/o ID Ours w/ ID GT
Figure 15: Ablation experiments about arcface identity embedding (ID). The additional identity embedding can greatly reduce the false illusion of identity information in the recovery results.
4.5 Limitations and Discussion
Although this represents an initial foray into attribute text-guided face image restoration, the flexibility of its text input is somewhat constrained due to the nature of the training samples. The model struggles to fully comprehend freely composed attribute description sentences, tending instead to rely on attribute labels embedded within fixed template text prompts, which limits its applicability in broader contexts. Furthermore, when users input attribute labels unseen during training, these do not effectively guide the recovery process. These limitations highlight the importance and necessity of utilizing high-quality data on a larger scale.
Moreover, we believe that allowing excessive free facial attribute control in FR tasks is undesirable due to the potential risk of model abuse. Our model can still recover well without attribute text prompts, thus incorporating additional multimodal information marks our initial effort to enhance recovery performance without increasing task complexity. Regarding the reference image, we disregard its potential to ”edit” the restored results. As with previous studies, we aim for the reference image to further enhance the facial details of the restored face image. In summary, we remain committed to the core objective of the restoration task, ensuring that the output remains faithful to the low-quality input. Moreover, given that individuals typically have multiple reference images, leveraging a broader range of reference information may result in more precise facial detail restoration. Additionally, we acknowledge that high-quality reference images may not always be available, thus the potential of utilizing multiple low-quality reference images for co-guidance is left for future investigation.
5 Conclusion
We introduce MGFR as a pioneering method in real-world face restoration, at the cutting edge of face image restoration technology, capable of using multi-modal information for guidance to achieve realistic visual effects. Simultaneously, MGFR extends the possibilities of face restoration by controlling text prompts with attributes. The proposed Reface-HQ dataset also offers significant potential for advancing the development of face restoration models based on reference images. As the first multi-modal face image restoration model, MGFR establishes a new benchmark for future technological advancements.
Appendix
Appendix A Reface-HQ Dataset
Dataset Number of ID Image Size Synthesized
CASIA-WebFace 10575 494414 256$\times$256
Celeba 10,177 202599 178$\times$218
IDiff-Face - - 128$\times$128
VggFace2 - HQ 1200 24000 512$\times$512
CelebRef-HQ 1000 10000 512$\times$512
Reface-HQ 5250 23500 512$\times$512
Table 5: Datasets Comparison.
The bulk of prior reference-based face restoration methodologies commonly focus on training and testing with $256\times 256$ images. This is primarily due to the limitations of existing datasets such as CelebA Liu et al. (2015), VggFace2 Cao et al. (2018), and CASIA-WebFace Yi et al. (2014), which offer reference images mainly for face or attribute recognition but do not include high-quality images suitable for training at higher resolutions, like $512\times 512$ or $1024\times 1024$, thereby limiting their practical applications. Additionally, high-definition datasets recently introduced, such as CelebRef-HQ Li et al. (2022) and VggFace2-HQ, face challenges in maximizing the potential of models due to their limited number of images and narrow range of identities.
To address this challenge, we have created a new real-world dataset named Reface-HQ, as shown in Figure 16. The Reface-HQ dataset encompasses high-definition facial images of celebrities, which have been collected from the Internet. Initially, images with inadequate resolution (minimum 512), low quality and outliers lacking facial features were eliminated. Subsequently, identities represented by fewer than two images were excluded, and face image crop alignment was conducted. Each identity was also manually inspected to eliminate discrepancies in age and makeup. Additionally, to enhance the fairness and inclusiveness of the algorithm, we meticulously review the dataset to ensure it includes samples from all races and skin colors. We strive to ensure the diversity of the training data, thereby minimizing algorithmic bias and discrimination, and further enhancing the algorithm’s fairness and inclusiveness. In summary, Reface-HQ encompasses 5,250 identities, totaling 23500 images with a resolution of 512, subsequently partitioned into three segments: 4870 identities for the training set and 380 for the Reface-Test. The comparison of datasets available for special face restoration tasks is shown in Table 5. IDiff-Face Boutros et al. (2023) is a composite dataset with an indefinite number of images.
Figure 16: Demonstration of the Reface-HQ dataset.
A.1 Ablation Experiment
For diffusion models and adapter structures, both the quality and quantity of training data are critical factors affecting the model’s final performance. Table 6 presents the quantitative comparison results of our proposed model under various training data volumes. It is evident that the model’s performance significantly decreases with only 10,000 training data samples.
Real-SR(×8) SSIM PSNR LPIPS ManIQA ClipIQA MUSIQ
10K Training Samples 0.6254 23.46 0.2535 0.6388 0.7424 72.23
20K Training Samples 0.6248 23.10 0.2688 0.6535 0.8147 75.51
Table 6: Ablation Experiment about training.
Appendix B Attribute Prompt
Row 1 Row 2 Row 3 Row 4
Black Hair Blond Hair Blurry Brown Hair
- Eyeglasses Gray Hair Heavy Makeup
Mouth Slightly Open Mustache Big Eyes No Beard
Receding Hairline Sideburns Smiling Straight Hair
Wearing Earrings Wearing Hat Male Wearing Necklace
Big Nose - Wearing Lipstick Young
Wavy Hair Big Lips Bald Bangs
Table 7: Face Attribute.
This section provides a supplementary note on the attribute text prompts utilized in MGFR. For the training data, attribute labels are first extracted from the FFHQ or Reface-HQ dataset’s face images using a facial attribute classifier. The 28 types of attributes included are listed in Table 7, while labels with binomial characteristics (such as Male and Female, no beard and beard, etc.) are not repetitively shown. Regarding the classification threshold, attributes with a probability greater than 0.6 are considered positive, those with a probability less than 0.4 as negative, and the rest as uncertain in describing facial features. LLM is utilized to embed the attribute labels into a descriptive sentence template, thereby enhancing the model’s understanding. To augment the model’s grasp of negative attribute descriptions, two sentences of prompt text are provided for each image, as illustrated in Figure 17. Both descriptions offer a positive portrayal of the face, with Prompt B specifically focusing on the negative attributes.
In the inference stage, following the approach detailed in Section 3.3, we apply positive attribute prompts ($pos$), negative quality prompts ($nq$), and negative attribute prompts ($na$) in each iteration. For example, in restoring a LQ image, if it is assumed to contain attributes like ’smiling, man, black hair, eyeglasses,’ the corresponding text for image restoration can be generated as follows:
  • Positive Prompt: A high quality, high resolution, realistic and extremely detailed image in the description of a smiling man who has black hair and eyeglasses.
  • Negative Attribute Prompt: A high quality, high resolution, realistic and extremely detailed image not in the description of a smiling man who has black hair and eyeglasses.
  • Negative Quality Prompt: A low quality, low resolution, over smooth and deformation image.
The underlying premise is to prevent our model from generating low-quality images and images with mismatched facial attributes. Extensive experiments demonstrate the effectiveness of our proposed attribute prompts.
Figure 17: Attribute prompts composition in training.
Appendix C More Qualitative Comparisons For Our Model without Reference Images.
WebPhoto-Test ManIQA ClipIQA MUSIQ
DR2 0.4868 0.6184 64.36
DiffBIR 0.4068 0.6858 55.73
Ours 0.5901 0.8397 72.52
Table 8: Quantitative comparison with other diffusion model-based methods on real-world degradations in WebPhoto-Test.
This section presents qualitative comparisons experimental results of our model without reference images, focusing on attribute text-guided face recovery. Importantly, for a fair comparison, the attribute during the inference phase are derived from the LQ input, which means the model’s maximum potential is not fully realized. We assert that in practical scenarios, users will be able to supply more precise attribute text for enhanced recovery guidance. Although, our model demonstrates the most superior visual effects and details when compared to other state-of-the-art methods.
Figure 18 and Figure 19 display the qualitative comparison results of our model against other advanced models under conditions of mild and moderate degradation of LQ input, respectively. It is evident that the previous methods exhibit severe facial illusion, whereas our model attains the best visual outcomes. Notably, as shown in Figure 20 and Figure 21, our model demonstrates a remarkable ability to recover severely degraded input images with high quality and fidelity. Finally, Figure 22 shows the effect of restoration on real-world LQ inputs and Table 8 presents the quantitative comparison results between our model and the principal comparison methods using real-world LQ inputs from the WebPhoto-Test dataset.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LR PSFRGAN GPEN DR2 VQFR DiffBIR CodeFormer Ours w/o Ref. GT
Figure 18: More qualitative comparisons for our text-guided baseline model on synthetic dataset under mild degradation in CelebA-Test dataset. Zoom in for best view.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LR GPEN DR2 VQFR DiffBIR CodeFormer Ours w/o Ref. GT
Figure 19: More qualitative comparisons for our text-guided baseline model on synthetic dataset under moderate degradation in CelebA-Test dataset. Zoom in for best view.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LR GPEN DR2 VQFR DiffBIR CodeFormer Ours w/o Ref. GT
Figure 20: More qualitative comparisons for our text-guided baseline model on synthetic dataset under severe degradation in CelebA-Test dataset. Zoom in for best view.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LR ASFFNet MDMNet DR2 CodeFormer DiffBIR Ours w/o Ref. GT
Figure 21: More qualitative comparisons for our text-guided baseline model on synthetic dataset under severe degradation in Reface-Test dataset. Zoom in for best view.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LR DR2 CodeFormer DiffBIR Ours w/o Reference
Figure 22: More qualitative comparisons for our model without reference images on real world images. Zoom in for best view.
C.1 Additional Supplement to Table 1
Due to space limitation, for the quantitative comparison results of our model without reference images, Table 1 does not show the numerical values of PSNR and SSIM, and we supplement them in Table 9, Table 10 and Table 11.
Real-SR($\times$4)
Method LPIPS PSNR SSIM ManIQA ClipIQA MUSIQ
PSFRGAN 0.2938 23.72 0.6522 0.5927 0.5702 73.39
GPEN 0.2828 24.78 0.7056 0.6596 0.6430 69.25
VQFR 0.2951 23.81 0.6878 0.2875 0.2490 62.95
CodeFormer 0.2927 24.56 0.6809 0.5803 0.5179 75.47
DR2 0.3264 23.74 0.6827 0.5749 0.4441 63.43
DiffBIR 0.2611 24.49 0.6778 0.6068 0.7681 74.27
MGFR(Ours) 0.2925 23.25 0.6104 0.6854 0.8244 76.22
Table 9: Quantitative Comparison in CelebA-Test. Results in red and blue signify the highest and second highest, respectively. The $\downarrow$ indicates metrics whereby lower values constitute improved outcomes, with higher values preferred for all other metrics.
Real-SR($\times$8)
Method LPIPS PSNR SSIM ManIQA ClipIQA MUSIQ
PSFRGAN 0.3315 22.85 0.6232 0.6015 0.5956 73.08
GPEN 0.3217 23.88 0.6822 0.6754 0.6299 68.63
VQFR 0.3277 23.16 0.6683 0.4163 0.2363 61.92
CodeFormer 0.3193 21.81 0.5799 0.5970 0.6235 75.09
DR2 0.3580 23.26 0.6725 0.5246 0.4494 59.46
DiffBIR 0.3017 23.47 0.6442 0.6058 0.7439 73.87
MGFR(Ours) 0.3227 22.34 0.5904 0.6776 0.8083 75.94
Table 10: Quantitative Comparison in CelebA-Test. Results in red and blue signify the highest and second highest, respectively. The $\downarrow$ indicates metrics whereby lower values constitute improved outcomes, with higher values preferred for all other metrics.
Real-SR($\times$16)
Method LPIPS PSNR SSIM ManIQA ClipIQA MUSIQ
PSFRGAN 0.3788 21.27 0.5899 0.5739 0.6274 71.76
GPEN 0.3831 22.22 0.6541 0.6618 0.5897 66.61
VQFR 0.3761 21.72 0.6413 0.6513 0.2148 60.49
CodeFormer 0.3821 21.19 0.5717 0.5803 0.5877 70.85
DR2 0.3796 21.06 0.6225 0.5160 0.5035 70.31
DiffBIR 0.4238 21.21 0.5654 0.5361 0.7164 67.41
MGFR(Ours) 0.3760 20.54 0.5452 0.6729 0.7944 75.76
Table 11: Quantitative Comparison in CelebA-Test. Results in red and blue signify the highest and second highest, respectively. The $\downarrow$ indicates metrics whereby lower values constitute improved outcomes, with higher values preferred for all other metrics.
Appendix D Inference Time and Memory Consuming Analysis
Our method involves two stages of training, and it is clear that our proposed model requires substantial training resources. However, for diffusion-based models, this duration of training is relatively short. We believe the investment is worthwhile as the proposed MGFR model delivers strong results. Additionally, Table 12 displays the average time and memory consumption for model inference. It is important to emphasize that the CFG strategy can be applied to all LDM-based recovery models. Consequently, the results are displayed separately to indicate the use of the CFG strategy during the inference process. Clearly, when our model does not utilize the CFG strategy, time and memory consumption improve only marginally compared to DiffBIR. DR2 boasts the fastest inference time; however, its restoration results are unsatisfactory due to the lack of an additional controller structure. Nevertheless, we assert that efficiency is not the primary focus of this work.
Method Average time (s) Memory consuming (M)
DiffBIR 5.1 11260
DR2 2.6 3144
Ours(w/o CFG) 6.9 15351
Ours (w/ CFG) 12.5 15351
Table 12: Inference Time and Memory Consuming compared with other diffusion model-based methods.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Input No Prompt ‘old, woman’ Input No Prompt ‘eyeglasses’
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Input No Prompt ‘earrings’ Input No Prompt ‘lipstick’
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Input No Prompt ‘mouth slightly open’ Input No Prompt ‘big eyes’
Figure 23: Influences of attribute prompts.
Appendix E Controlling with Attributes Prompts
E.1 Controlling Restoration
Our model facilitates guidance through user-defined attribute prompts during testing. Figure 23 exemplifies this with a demonstration of attribute prompt-controlled recovery. Notably, ’No Prompt’ refers to the initial prompt input, ’A high quality, high resolution, realistic, and extremely detailed image.’ As illustrates, users can employ prompts like ’old’ to define the approximate age in the restored image, or ’eyeglasses’ and ’earrings’ to add accessories to the image. Furthermore, users can provide additional attribute prompts to refine unsatisfactory results. For instance, ’lipstick’ can be used to add lipstick, or ’mouth slightly open’ to adjust the mouth’s appearance. More significantly, severe illusions, particularly in the eye area, are common in previous methods due to insufficient information in LQ inputs. This observation underscores the importance of attribute prompts in our method, as using ’big eyes’ leads to more realistic eye effects. Therefore, we posit that attribute text holds potential as a versatile tool for controlling face recovery.
E.2 Sensitivity Analysis
Moreover, as depicted in Figure 24 case 1 and case 2, with increasing levels of degradation, the model’s reliance on attribute prompts for control becomes more apparent, leading to greater flexibility. This observation, a logical experimental outcome, confirms the model’s fidelity to LQ inputs during recovery. Specifically, attribute prompts that starkly contradict the LQ input do not influence the effect, which aligns with our expectations. The primary function of attribute labels, we contend, is to facilitate more efficient and effective image restoration, rather than to focus on image editing and control. This is intrinsic to the core objective of real-world face restoration. In our method, all attribute labels listed in Table 7, including ’black hair’, ’brown hair’, and others, do not possess the ability to control recovery but rather aid the model in interpreting the LQ input. These insights robustly underscore the effectiveness of our approach.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Input ($\times$ 4) No Prompt ‘no glasses’ Input ($\times$ 8) No Prompt ‘no glasses’ Input No Prompt ‘blue eyes’
Figure 24: We investigate the following options for attribute prompt control. First of all, the model becomes increasingly dependent on attribute prompt as input deterioration increases (case 1 & case 2). Second, the input attribute tag does not have a control role if it is not present in Table 7 (case 3).
Appendix F User Study
Currently, the relevance and efficacy of metrics such as PSNR, SSIM, and LPIPS require evaluation. In this study, a User study was conducted as an alternative metric for assessing image restoration quality. The study concentrated on two primary questions: (1) How does our model without reference images perform in terms of restoring image quality versus reducing facial illusions compared to previous methods? (2) Does the addition of reference image and identity information in guiding restoration result in images that are closer to the Ground Truth compared to the model without reference images? Two sets of questionnaires were prepared, and the study was conducted with 50 participants. Participants were presented with random, anonymous options for their selection. For question (1), our model was compared with DiffBIR Lin et al. (2023), VQFR Gu et al. (2022), and CodeFormer Zhou et al. (2022), focusing on selecting images with better quality and fewer hallucinations, without providing Ground Truth images. This comparison involved 50 sets of images. For question (2), a self-comparison approach was adopted. Specifically, ground truth images were provided, and participants were asked to choose between restoration results with and without reference images, assessing them based on their proximity and realism to the ground truth. In this experiment, 50 pairs of synthetically degraded images were compared.
Subsequently, the first part of the user study, focusing on the improvement of our model in terms of image quality and the reduction of facial illusion, is discussed. The results and detailed information of this study segment are presented in Figure 25 and Figure 26. It was observed that the majority of the 50 participants favored our model for its superior image quality and minimal facial illusions. Reflecting on the recovery results of the advanced method CodeFormer, illustrated in Appendix C, it is noted that while CodeFormer achieves relatively good quality in restored images, considerable facial illusions persist, particularly around the mouth and eyes. In contrast, our method consistently produces high-quality, realistic facial images with minimal facial illusion. These findings underscore the our model’s capability to reduce illusion and enhance image quality through negative prompts. Specifically, supported by the diffusion model and LR control adapter, our model is adept at generating realistic high-quality restorations influenced by negative prompts, and it effectively minimizes facial illusions by utilizing an optimal amount of attribute prompts. The synergy of these elements paves the way for further exploration in MGFR.
It is noteworthy that our two-part User study also corresponds to the two-stage development process of the MGFR model. For the second part, the first User study has demonstrated the superior performance of our model, as shown in the figure. Participants generally agreed that adding a guide to the reference image would further achieve superior visual effects.
Figure 25: Results and question details of user study.
Figure 26: Results and question details of user study.
Appendix G Ablation Study for Negative Prompt
For negative prompts, we introduce two hyperparameters, $\lambda_{na}$ and $\lambda_{nq}$. However, we find that the changes of the two values tend to have the same effect on the restored images. Thus, we keep $\lambda_{na}=0.5$ and $\lambda_{nq}=0.5$ during reasoning. Here we will represent the values of $\lambda_{na}$ and $\lambda_{nq}$ with $\lambda$ to show the qualitative comparison results under different hyperparameters in Figure 27.
Appendix H Impact Statements
Controlled generation technology, as a pivotal innovation in the field of diffusion models, exerts a significant impact across multiple sectors of society. In the creative industries, it enables artists and designers to realize complex visions with unprecedented precision and flexibility, fostering innovation in digital art, design, and multimedia content creation. In commercial applications, controlled generation technology enhances marketing strategies by offering more targeted and dynamic advertising visuals, effectively engaging consumers. Additionally, its influence extends to education and training, where it can revolutionize teaching methods and materials, especially in visually-dependent disciplines, by generating customized educational content and simulations.
The work presented in this paper aims to advance machine learning and computer vision. This method can provide the public with better face processing effects and has greater social value. However, the technique is designed to process facial information, inevitably involving facial attributes such as race and privacy risks. We are aware of these risks. Our research uses publicly available data and images accompanied by captions. We are also wary of potentially discriminatory attribute descriptions in our research. Our method also provides control over face restoration, which reduces the possibility of our method outputting harmful information.
Appendix I More Qualitative Comparisons For MGFR Model
Figure 28 displays the qualitative comparison results between the proposed MGFR model and other advanced methods. The “w/o Reference Image” represents the restoration results of our model after initial training. The use of the negative intuition strategy and attribute prompts significantly reduces the false illusions in face images and substantially enhances overall quality. Subsequently, the inclusion of additional multi-modal information, such as reference images and identity information, can achieve superior visual effects.
Figure 27: Influence of hyperparameters on recovery effect in CFG. The smaller $\lambda$ does not get a clear recovery result and the huge $\lambda$ causes the recovered image to be over sharp.
Figure 28: More qualitative comparisons for MGFR with reference image and ID guidance on synthetic dataset in Reface-Test dataset. Zoom in for best view.